OpenStudy (anonymous):

-5x-8y=1 what's that in slope intercept form

OpenStudy (anonymous):

If you meant slope intercept form this is it. -5x-8y=1 Add 5x on both sides +5x +5x -8y=5x+1 Divide both sides by -8 /-8 /-8 /-8 y=-5/8x-1/8 Hope I helped =]

OpenStudy (anonymous):

Thanks it helped a lot, jus got confused on the dividing part

OpenStudy (anonymous):

You divide by -8 because your trying to get y by itself, and what you do on one side of the equation must be done on the other side. Also because it's slope intercept form you don't actually have to divide unless you get a whole number.